OpenStudy (anonymous):

if you start at the number 15 and there are three missing numbers and the fourth number ids 355 , what is the pattern , and the three missing numbers.

OpenStudy (anonymous):

please help

OpenStudy (asnaseer):

if the numbers are in an arithmetic progression, then there will be a constant difference between each one - call that 'd'. so the number we have form this sequence: 15, 15+d, 15+2d, 15+3d, 15+4d = 355 use the last expression to work out the value for 'd' which will then allow you to work out the three missing numbers.

OpenStudy (anonymous):

please can you just tell me the pattern and i'l do the rest myself.

OpenStudy (anonymous):

please @asnaseer

OpenStudy (asnaseer):

I've shown you what to do above ^^

OpenStudy (asnaseer):

solve the following to find 'd':\[15+4d=355\]

OpenStudy (asnaseer):

'd' represents the difference between each term of sequence - that is your 'pattern'

OpenStudy (anonymous):

thank you very much

OpenStudy (asnaseer):

yw

OpenStudy (asnaseer):

I see you haven't closed this question yet - are you still stuck? I thought you understood now how to solve it?

OpenStudy (anonymous):

i don't get it

OpenStudy (asnaseer):

ok - let me try to explain step-by-step...

OpenStudy (asnaseer):

you are given the first number: 15 and you are given the last number: 355

OpenStudy (asnaseer):

you are also told that there are three numbers inbetween

OpenStudy (anonymous):

k

OpenStudy (asnaseer):

so lets say the difference between each of these numbers remains the same - i.e. they form an arithmetic sequence - do you know what that is?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

an arithmetic sequence is one where you have a sequence of numbers where the difference between each successive number is the same. e.g. the sequence: 2, 7, 12, 17, ... is an arithmetic sequence because you just add 5 each time to get from one term to the next.

OpenStudy (asnaseer):

make sense?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

later you should check some web sites to become more familiar with this if you want, e.g. try this one: http://www.basic-mathematics.com/arithmetic-sequence.html

OpenStudy (asnaseer):

now, in your problem, you are given the first and last numbers of this sequence and you are told that there are 3 missing numbers in between

OpenStudy (asnaseer):

now what I did was use a variable 'd' to represent the difference between each term.

OpenStudy (asnaseer):

we need to work out what this common difference is

OpenStudy (asnaseer):

understand?

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

so, if the difference between each of our terms is 'd', then we know we can write the sequence as: 15, <--- 1st term 15 + d, <--- 2nd term 15 + d + d , <--- 3rd term 15 + d + d + d, <--- 4th term 15 + d + d + d + d <--- 5th term

OpenStudy (asnaseer):

because each time we add on another 'd' to get to the next term - understand?

OpenStudy (asnaseer):

if it helps, we can look at the example sequence we saw above: 2, 7, 12, 17, ... this we said had a difference of 5 between each term, so it can be written as: 2, 2 + 5 (=7), 2 + 5 + 5 (=12), 2 + 5 + 5 + 5 (=17), ...

OpenStudy (asnaseer):

do you understand so far?

OpenStudy (asnaseer):

hello? if you don't reply then it is very difficult to help...

OpenStudy (anonymous):

yeah so sorry about that

OpenStudy (anonymous):

i really get it now .

OpenStudy (asnaseer):

ok - so no more explanation required?

OpenStudy (anonymous):

yeah